OpenStudy (anonymous):

th first flower bed that chris made was 4 1/4 bags of barkdust. the flowerbed he is now making is 1 1/3 as large as the first. how much barkdust will chris need?

OpenStudy (anonymous):

4 + 1/4 = 17/4 1 + 1/3 = 4/3 17/4*4/3 = 17/3

OpenStudy (anonymous):

hey do you mind helping me with my last question about the oil??((: